OpenStudy (anonymous):

Can someone please help? I'm really bad at mouth.. 1. Which method would be the simplest way to solve the system? 7x+5y=19 -7x-2y=-16 **Graphing **Substitution **Elimination **Disributive

OpenStudy (anonymous):

Haha I meant math.. Lol could you possibly help with a few more??

OpenStudy (anonymous):

elimination because we can remove the 7x and -7x and simply to one equation in terms of y to solve for y

OpenStudy (anonymous):

What is the solution of the system? 7x+5y= 19 -7x-2y=-16 **(2,1) **(1,2) **(7,5) ** (-7,-2)

OpenStudy (anonymous):

Its fine mechanical animal :)

OpenStudy (anonymous):

3.) Which method would be the simplest way to solve the system? y=1/2x 2x+3y=28 **graphing **substitution **Elimination **Distributive 4.) What is the solution of the system? y= -3x 3x+2y=6 ** (2,6) ** (-1,3) ** (-2,6) ** (1.5, -4.5) 5.)How many solutions does this system have? -3x+6y=10 -3x+6y= -4 ** none ** one **infinite **two 6.) In which quadrant does the solution of the system fall? y=x-1 y=-3x-5 ** I **II **III **IV 7.) In which quadrant does the solution of the system fall? x+y= 4 2x-y=2 **I **II **III **IV 8.) The cost of four sccarves and six hats is $52.00 The cost of two hats is $1.00 more than the cost of one scarf. What is the cost of one scarf? **$4 **$5 **$6 **$7

OpenStudy (anonymous):

go on mathway it gives u the answers

OpenStudy (anonymous):

Mathway??

OpenStudy (anonymous):

I dont want just the answers, I wanna know too

OpenStudy (anonymous):

Thank you, can you help with any of the others